Two sets A and B are such that n(universal set) = 90, n(A) = 40 and n(B) = 68.Find
a.The smallest possible value of n(A U B),
b.The greatest possible value of n(A U B),
c.The greatest possible value of n(A n B').

Answer by ikleyn(53937)

.
Two sets A and B are such that n(universal set) = 90, n(A) = 40 and n(B) = 68.Find
a.The smallest possible value of n(A U B),
b.The greatest possible value of n(A U B),
c.The greatest possible value of n(A n B').
~~~~~~~~~~~~~~

(a)  The smallest possible value of n(A U B) is 68,  when the set A is a subset of the set B.



(b)  The greatest possible value of n(A U B) is 90, when the sets A and B are maximally disjoint (= as disjoint as possible).



(c)  n(B') = 90 - 68 = 22.

     Therefore, the greatest possible value of n(A n B') is 40 + 22 = 62, when the sets A and B' are disjoint.

Solved.     //     All questions are answered.